OpenStudy (anonymous):

sketch the graph of f f(x)= x+ lx-1l/x-1

OpenStudy (turingtest):

\[f(x)=x+\frac{|x-1|}x-1\]?

OpenStudy (anonymous):

\[x+\frac{ \left| x-1 \right| }{ x-1 }\]

OpenStudy (turingtest):

break this up into intervals....

OpenStudy (turingtest):

There are two cases\[x-1<0\implies x<1\implies{|x-1|\over x-1}=-1\]\[x-1>0\implies x>1\implies{|x-1|\over x-1}=1\]so we can make a step function\[f(x)=\left\{\begin{array}yx-1~~~~\text{ for }~~~~x<1\\x+1~~~~\text{ for }~~~~x>1\end{array}\right.\]

OpenStudy (turingtest):

at \(x=1\) f(x) is undefined

OpenStudy (turingtest):

you should be able to graph that step function

OpenStudy (anonymous):

alright thanks! :)

OpenStudy (turingtest):

welcome!